sqrt(4x-5) = sqrt(x+9)

sqrt(x+49) + 7=x


Each time that I tried to solve these two (seperate) problems....kept getting an answer that does not make any sense....I think that I am doing something wrong but not sure where I am screwing it up. Can someone show me what I am doing wrong?- thank you.

Answer by stanbon(75887)

sqrt(4x-5) = sqrt(x+9)
---
Square both sides to get:
4x-4 = x+9
3x = 13
x = 13/3
--------------------------------
sqrt(x+49) + 7=x
sqrt(x+49) = x-7
Square both sides to get:
x+49 = x^2-14x+49
x^2 -15x = 0
x(x-15) = 0
x = 0 or x = 15
------------
Checking for extraneous roots:
Can x = 0 ?
sqrt(0+49)+7 = 0
7+7 = 0
false
---
Can x = 15 ?
sqrt(15+49) +7 = 15
sqrt(64)+7 = 15
8+7 = 15
true
========
Only solution: x = 15
==============================
Cheers,
stan H.
--------------------